OpenStudy (ttp):

A police officer was sent to investigate a car accident. He knows the speed of the car, s, can be measured with the equation s = at, where a is the rate of acceleration and t is the time. Which equation can he use to solve for the rate of acceleration? a = s – t a = t – s a = s ÷ t a = st

OpenStudy (ttp):

@esshotwired

OpenStudy (ttp):

@e.cociuba

OpenStudy (e.cociuba):

a = s ÷ t Take the information that the paragraph provides you, apart. You will come out with your answer as "a" equals "s" divided by "t".

OpenStudy (ttp):

A?

OpenStudy (esshotwired):

I agree with @e.cociuba but you can do it a different way also. You can take the equation s=at and divide both sides by s so it would be \[\frac{ s }{ at }\] then multiply both sides by a to get \[a=s \div t\]

OpenStudy (esshotwired):

it isn't A

OpenStudy (e.cociuba):

"where a is the rate of acceleration...." So you know by this that it's: a/at Multiply both sides to get your answer: a= s/t

OpenStudy (e.cociuba):

Yup, agreed @esshotwired :D

OpenStudy (ttp):

It's C!

OpenStudy (e.cociuba):

Correct! Good job! :)

OpenStudy (ttp):

Thank you ! :)

OpenStudy (e.cociuba):

Np! :)